OpenStudy (anonymous):

Solve the equation for x: √(12 – x) = 11.

OpenStudy (anonymous):

@Squirrels

OpenStudy (anonymous):

square both sides and start with \[12-x=121\]

OpenStudy (anonymous):

then it takes two steps to solve for \(x\) you good from there?

OpenStudy (anonymous):

where did yo get 121?

OpenStudy (anonymous):

i squared both sides and \(11^2=121\)

OpenStudy (anonymous):

and than 121/12?

OpenStudy (anonymous):

oh no lets go slow

OpenStudy (squirrels):

You subtract 12 from both sides at that point since you want to get x by itself

OpenStudy (anonymous):

\[\sqrt{12-x}=11\] square both sides, get \[12-x=121\] subtract \(12\) get \[-x=109\]

OpenStudy (anonymous):

then change the signs and your final answer is \(x=-109\)

OpenStudy (anonymous):

Oh..

OpenStudy (squirrels):

Do you still get it? or do you need it to be clarified?

OpenStudy (squirrels):

Ill explain it easier if you still need help

OpenStudy (anonymous):

Yeah if you could. I don't see why the number was positive and then he switched to negative.

OpenStudy (squirrels):

Ok do you want me to start from the beginning or just that part?

OpenStudy (anonymous):

beginning is fine.

OpenStudy (squirrels):

Ok so the equation is √(12 – x) = 11. You see that square root? -> √(12 – x) = 11. Well we want to get rid of that to get x by itself. So what is the opposite of square root? ^2

OpenStudy (squirrels):

Oh wow hold on this site is messed up

OpenStudy (anonymous):

It is.

OpenStudy (squirrels):

Sorry i mean here http://www.twiddla.com/1817159